OpenStudy (anonymous):

lets try this again i need help step by step for this equation square root of 4x plus 5 minus the square root of x minus one equals 3

OpenStudy (anonymous):

\[\sqrt{4x + 5} - \sqrt{x - 1} = 3\] \[\sqrt{4x + 5} = 3 + \sqrt{x - 1}\] \[4x + 5 = ( 3 + \sqrt{x - 1} )^2\] \[4x + 5 = 9 + 6\sqrt{x - 1} + x -1\] \[4x + 5 = 8 + 6\sqrt{x - 1} + x \] \[4x + 5 -x -8 = 6\sqrt{x - 1} \] \[3x - 3 = 6\sqrt{x - 1} \] \[ (3x - 3)^2 = (6)^2(x - 1) \] \[ 9x^2 - 18x + 9 = (36)(x - 1) \] \[ 9x^2 - 18x + 9 = 36x - 36 \] \[ 9x^2 - 18x + 9 -36x + 36 = 0 \] \[ 9x^2 - 54x + 45 = 0 \] \[ (3x - 3)(3x - 15 ) = 0 \] \[ 3x - 3 = 0 ; 3x - 15 = 0 \]

OpenStudy (anonymous):

\[ 3x - 3 = 0 \] \[ 3x = 3 \] \[ x = 1 \]

OpenStudy (anonymous):

\[ 3x - 15 = 0 \] \[ 3x = 15 \] \[ x = 5 \]

OpenStudy (anonymous):

Is that okay??

OpenStudy (anonymous):

hang on one sec

OpenStudy (anonymous):

omg thank you so much i have been stguck on this prob for 2 days

OpenStudy (anonymous):

What grade are you? And BTW, don't say OMG because you're putting the name of God in vain. (If that's what you ment my that abbreviation) :))

OpenStudy (anonymous):

*meant

OpenStudy (anonymous):

*by. LOL Typo.

OpenStudy (anonymous):

math 100 and no its oh my goodness i dont use God in in in vain vocab only in postive speaking

OpenStudy (anonymous):

and i know you r probably saying wow u r in only math 100 and u needed help but it has been 14 years since i have taken a math course thank you for your help

OpenStudy (anonymous):

what's a math 100? LOL. There's no Math 100 here in the Philippines. XD

OpenStudy (anonymous):

intermediate algrebra for college math befor calculus

OpenStudy (anonymous):

Oh. College math. So, you're a college student?